OpenStudy (anonymous):

\[\sum_{n=1}^{\infty} \frac{10^{n}x^{n}}{n^{3}}\]

OpenStudy (anonymous):

I have the radius of convergence as 1 and \[\lim_{n \rightarrow \infty} \left|x\right| \frac{10n^{3}}{(n+1)^{3}}\]

OpenStudy (anonymous):

i still need to find the interval of convergence

OpenStudy (anonymous):

quick question...wouldn't R always have to be <1 to be convergent?

OpenStudy (experimentx):

try using limit test

OpenStudy (anonymous):

isn't that what I did...I guess I didn't complete it

OpenStudy (anonymous):

\[\lim_{n \rightarrow \infty} \left|x\right| \frac{10}{(1+1/n)^{3}}\]

OpenStudy (experimentx):

Hmm ... that gives, radius of convergence is 1/10

OpenStudy (anonymous):

why 1/10 and not 10

OpenStudy (experimentx):

if x were 10 ... then just imagine what would happen to your series ... plug this into wolfram alpha and see ... the series will be divergent.

OpenStudy (anonymous):

ok...so now the interval is (1/10,1/10]

OpenStudy (anonymous):

im sorry

OpenStudy (anonymous):

(-1/10,1/10]

OpenStudy (anonymous):

haha

OpenStudy (experimentx):

no ... that's your radius of convergence.

OpenStudy (anonymous):

sigh...ok lets see

OpenStudy (anonymous):

does it have to do with numerator and denominator with the n terms (n+1)^3 and n^3

OpenStudy (anonymous):

no i'll take that back

OpenStudy (anonymous):

ok so i test the end points of R to see if it converges

OpenStudy (experimentx):

wait ... i'll be right back.

OpenStudy (anonymous):

- infty to +1/10 because it doesnt matter how small n becomes

OpenStudy (anonymous):

k

OpenStudy (experimentx):

it seems that your series converges for |x| < 1/10 so your interval of convergence is -1/10 < x < 1/10

OpenStudy (anonymous):

how is that different from (-1/10,1/10]

OpenStudy (anonymous):

or (-1/10,1/10)

OpenStudy (experimentx):

1/10 is closed on right hand side ..

OpenStudy (anonymous):

(-1/10,1/10)

OpenStudy (experimentx):

let's do it from the start.